OpenStudy (anonymous):

A cargo plane flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220 mph. The average speed on the way back was 200 mph. How many hours did the trip there take?

OpenStudy (anonymous):

yeah its a hard one

OpenStudy (anonymous):

the answer is 10 hours but i don't know how to work that problem

OpenStudy (anonymous):

i just have the answers and quizzing myself

OpenStudy (anonymous):

We know: Speed(to) = 220 Time(to) = x hours Speed(back) = 200 Time(back) = x+1 hours Speed = Distance / Time => Distance = Speed * Time Distance is same so equations for distance(to) and distance(back) can be equated so that: Speed(to) * Time(to) = Speed(back) * Time(back) 220*x = 200*(x+1) 220x = 200x + 200 220x - 200x = 200 20x = 200 x = 10 Therefore: Time(to) = x = 10 hours and Time(back) = x+1 = 10+1 = 11 hours.

OpenStudy (anonymous):

would d=220(x-1) have been sufficient?

OpenStudy (anonymous):

I wouldn't think so, you'd have to get the numerical answer as it is possible.

OpenStudy (anonymous):

Do you understand how to get it from what I've put?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

yeah i do thank you that was brilliant!

OpenStudy (anonymous):

No probs mate!